# Generalized Pinching-Antenna Systems: A Tutorial on Principles, Design Strategies, and Future Directions

### v1

@@ II-A1 System Architecture and Physical Implementation
A typical pinching-antenna system has three components: a base station (BS) acting as a central processing unit (CPU) to perform signal processing, dielectric waveguides employed as transmission media, and pinching antennas (small dielectric elements) for dynamic antenna deployment. As illustrated in Fig. 1(a), the dielectric waveguide is typically connected to the BS through a wired medium. Pinching antennas are strategically placed or dynamically activated along the waveguide, enabling on-demand radiation of electromagnetic energy into free space to form localized communication areas. This unique reconfigurability facilitates rapid deployment and flexible coverage, substantially enhancing the operational flexibility compared with conventional-antenna systems.
Figure 3: A schematic illustration of signal propagation and channel model for a single-input single-output (SISO) pinching-antenna system.
@@ II-A2 Signal Propagation and Channel Model
To illustrate the signal propagation features and derive the corresponding channel model of pinching-antenna systems, we consider a basic SISO scenario. In this case, a single pinching antenna mounted on a dielectric waveguide serves one single-antenna user in the downlink, as depicted in Fig. 3. Without loss of generality, we define the coordinate of the waveguide’s feed point as $\bm{\psi}_{0}=[0,0,d_{v}]$, where $d_{v}$ is the height of the waveguide. The coordinates of the pinching antenna and the user are denoted by $\bm{\widetilde{\psi}}^{\textrm{Pin}}=[\tilde{x},0,d_{v}]$ and $\bm{\psi}=[\bar{x},\bar{y},0]$, respectively.
@@
  • Stage 1 (In-waveguide propagation): The transmit signal $s$ is first fed into the waveguide at the feed point, after which it propagates through the waveguide’s core to the position of the pinching antenna. Due to the inherent characteristics of the waveguide, the signal experiences both attenuation and phase shifting. Specifically, the signal reaching the pinching antenna can be represented as
@@
    where $h^{\textrm{iw}}$ denotes the in-waveguide propagation channel gain, $n_{\textrm{eff}}$ represents the effective refractive index of the dielectric waveguide, $\lambda$ is the free-space wavelength, and $\alpha$ denotes the in-waveguide attenuation coefficient [^1] The in-waveguide attenuation coefficient depends on several factors, including the waveguide material, structure, and operating frequency. For example, in a typical dielectric waveguide, it is approximately $0.08$ dB/m [28, 29]. We include in-waveguide attenuation in the channel model for notational completeness; however, as will be discussed in Section III, its impact on system performance is generally negligible under mild and practically achievable system parameter settings..
@@
    where $h^{\textrm{fs}}$ denotes the free-space channel gain, $\eta=\frac{c^{2}}{(4\pi f_{c})^{2}}$ comes from the Friis equation and captures the frequency dependence of free-space propagation loss [31]. Here, $c$ denotes the speed of light and $f_{c}$ represents the carrier frequency.
By combining the channels of these two stages, the overall effective channel gain from the waveguide feed point to the user can be modeled as
@@
$=$ $\underbrace{\frac{\eta^{\frac{1}{2}}}{\|\bm{\psi}-\bm{\widetilde{\psi}}^{{\rm Pin}}\|}}_{\text{free-space path loss}}\underbrace{\exp(-\alpha\tilde{x})}_{\text{in-waveguide attenuation}}$
$\exp\bigg({-j\bigg(\underbrace{\tfrac{2\pi}{\lambda}\|\bm{\psi}-\bm{\widetilde{\psi}}^{{\rm Pin}}\|+\tfrac{2\pi}{\lambda_{g}}\tilde{x}}_{\text{overall phase shifts}}\bigg)}\bigg),$ (3)
where $\lambda_{g}=\frac{\lambda}{n{\textrm{eff}}}$ denotes the guided wavelength within the waveguide.
The channel model derived above highlights several distinctive properties of dielectric waveguide-based pinching-antenna systems. First, unlike the traditional antenna systems where the antenna positions are fixed, the flexibility to dynamically position pinching antennas along the waveguide enables fine-grained control over path loss. By strategically adjusting antenna placement, the system can reduce propagation distances and establish strong LoS connections, thereby improving the received signal strength. Second, the overall channel gain in pinching-antenna systems captures the phase shifts resulting from both in-waveguide and free-space propagation. This dual phase accumulation, which is absent in conventional wireless channels, means that the channel phase can also be reconfigured through the mechanical repositioning of the antenna or by exploiting the waveguide’s propagation characteristics. Such tunability is not available in conventional fixed-antenna systems. Altogether, these features provide pinching-antenna systems with better adaptability in channel shaping and user-centric link creation. At the same time, they introduce new dimensions of complexity, presenting new challenges and opportunities for the development of optimization algorithms for efficient pinching-antenna system designs.
@@ II-B LCX-Based Pinching Antennas
LCX, also known as radiating cable, provides a practical and user-centric means to realize flexible antenna systems, making it a natural extension of the pinching-antenna concept. As shown in Fig. 1(b), in LCX-based systems, the coaxial cable has periodic slots or apertures, which enable controlled and distributed electromagnetic energy leakage at specific locations [14]. This allows the cable to function as a long, linear antenna array that supports seamless wireless coverage—an approach widely adopted in underground tunnels, subway systems, and other challenging environments.
@@
TABLE III: Comparison of Representative Generalized Pinching-Antenna Systems
Pinching-Antenna System Type Signal Propagation Medium Propagation Mode In-Medium Signal Loss Operating Frequency Active Antenna Required Deployment Scenario Medium Cost
Dielectric-Waveguide-Based Pinching-Antenna System Dielectric waveguide Guided mode Low High No Indoor & Semi-Outdoor Medium
LCX-Based Pinching-Antenna System Coaxial cable with engineered slots Quasi-TEM mode Medium Low No Indoor & Outdoor Medium
Pinching-Inspired Antenna System Surface of conductor (e.g., metal pipes) Surface-wave mode High Low & High Yes Indoor & Outdoor Low
@@ II-C Pinching-Inspired Antenna Systems
The key idea of pinching antennas, realizing on-demand distributed radiation along conductive or engineered surfaces, can be used to design various pinching-antenna-inspired antenna array systems. In particular, signals are transmitted/received through two stages in both waveguide-based and LCX-based pinching-antenna systems. The first stage is essentially wired transmission. For example, in DOCOMO’s pinching-antenna systems, signals are passed through a dielectric waveguides. The second stage is wireless transmission. For example, in DOCOMO’s pinching-antenna system, signals are radiated by pinches and sent to the users wirelessly. Therefore, when designing pinching-inspired antenna systems, one question is whether there are other possible signal propagation media in addition to dielectric waveguides and LCX. Ideally, such media should be readily available in practical communication environments and incur a lower cost than dielectric waveguides.
One promising example for the design of pinching-inspired antenna systems is shown in Fig. 1(c), which is motivated by the following fact. In practical IoT scenarios, there are many metal wires and pipes, such as water and chimney pipes. Through surface-wave communications, these metal pipes can deliver signals to nearby users. Unlike the guided model used by dielectric-waveguide-based transmission, the signals are passed along a metal pipe in the so-called surface-wave mode, i.e., the signals are trapped and propagated on the surface of the pipe [33]. However, similar to dielectric-waveguide-based pinching antennas, the first phase of transmission is still wired propagation. Due to the low-cost nature of the medium being used, this new type of pinching-inspired antenna system can be realized with a lower cost than dielectric waveguide-based pinching-antenna systems. However, this new type of pinching-antenna system based on the surface-wave mode faces two challenges. First, signals propagated via the surface-wave mode may experience higher loss and interference than signals propagated via the guided mode of dielectric waveguides. The second challenge is how to flexibly realize the second stage of pinching-antenna systems, i.e., how to dynamically radiate signals to the user. As shown in Fig. 1(c), one approach is to use active antennas instead of passive antennas, as in waveguide/LCX-based systems. Consider an IoT communication scenario in which a user uses augmented reality (AR) or virtual reality (VR) for smart factory–related training. Conventionally, the serving access point is fixed at the center of the factory and cannot provide a strong LoS link to this user. With the pinching-inspired antenna system, the user can manually attach an active antenna to the surface where a strong LoS link is available and remove the antenna after the training session.
@@
Table III provides a comprehensive comparison of three representative generalized pinching-antenna systems, highlighting how each leverages unique physical and signal propagation characteristics to achieve flexible, on-demand wireless coverage. The “Propagation Mode” column describes the fundamental electromagnetic mode by which signals are transported within the medium. For dielectric-waveguide-based pinching-antenna systems, signals are conveyed using the guided mode, which confines the electromagnetic wave within the dielectric structure, resulting in low in-medium signal loss and enabling high-frequency operation. In contrast, LCX-based systems utilize a quasi-TEM (transverse electromagnetic) mode characteristic of coaxial cables with engineered slots; this mode provides controlled leakage for distributed coverage but incurs higher signal loss and is typically used at lower frequencies. Pinching-inspired antenna systems, built on metal pipes or radio stripes, exploit surface-wave modes in which the signals propagate along the surface of the medium and then radiate into free space at the desired position. These systems support both low- and high-frequency operation but experience higher in-medium attenuation. These systems also differ in their typical deployment scenarios. Specifically, dielectric-waveguide-based pinching-antenna systems are suitable for indoor settings and semi-outdoor environments, such as train stations, stadiums, or corridors that are not fully exposed to outdoor elements but are not entirely enclosed. LCX-based and pinching-inspired systems, due to their robust structural designs, are well suited for both indoor and outdoor deployments. Overall, this comparison underscores the versatility and adaptability of the generalized pinching-antenna framework, showing how diverse physical mechanisms and design choices enable tailored wireless solutions for a wide variety of coverage needs and environments.
@@ III-A Single-Waveguide Systems
Pinching-antenna systems having a single dielectric waveguide represent the most basic yet versatile architecture for channel reconfiguration in wireless networks. Depending on the number of pinching antennas deployed and activated, as well as the user configuration, the system can realize a broad range of functionalities and performance trade-offs, as shown in Fig. 4. In this subsection, we discuss two scenarios where a waveguide is equipped with a single pinching antenna and multiple pinching antennas, respectively. For each scenario, we consider both single-user and multiple-user transmission.
@@ III-A1 Single Pinching Antenna on the Waveguide
According to [36, Lemma 1], the globally optimal solution to problem (4) can be derived in closed-form and is given by
@@
From a data rate maximization perspective, the analysis in [36] reveals that ignoring waveguide attenuation is justified, if certain mild and practically achievable conditions on the system parameters, such as waveguide height, attenuation coefficient, and the size of the communication region, are satisfied. Specifically, according to [36, Proposition 1], the average data rate loss introduced by neglecting waveguide attenuation can be approximated as
@@
to ensure that the average rate loss from ignoring in-waveguide attenuation remains below a chosen threshold $\epsilon$. For example, if $d_{v}=10$ m, $\alpha=0.0092$ m-1 (corresponds to $0.08$ dB/m), and the target rate loss is $\epsilon=0.1$ bps/Hz, we have $D\leq 92.88$ m. This demonstrates that, for typical system parameters, the region over which the in-waveguide attenuation can be safely neglected is sufficiently large for most practical deployments. More recently, the impact of in-waveguide attenuation has been studied in a more general scenario with probabilistic LoS blockage in [37]. According to the analysis in [37, Proposition 1], the average data rate loss due to ignoring in-waveguide attenuation is given by
$\mathbb{E}_{{\bm{\psi}}}[\Delta R]$
$\approx\frac{\alpha^{2}}{\beta\ln 2}\left[1\!-\!\frac{2}{D\sqrt{\beta(1\!+\!\beta d_{v}^{2})}}\arctan\left(\frac{\sqrt{\beta}\,D/2}{\sqrt{1\!+\!\beta d_{v}^{2}}}\right)\right],$ (8)
where $\beta\leq 1$ is a small positive constant and denotes the blockage density parameter. It is typically chosen between $0.01$ and $0.1$ m-2 [38] and represents the density of obstacles in the propagation area. A higher $\beta$ value corresponds to a denser blockage environment, thereby reducing the probability of sustaining a LoS connection. The expression in (8) establishes how the average rate degradation is jointly influenced by key system parameters, including the in-waveguide attenuation coefficient $\alpha$, the blockage density parameter $\beta$, the waveguide deployment height $d_{v}$, and the communication area size $D$. Based on this characterization, several practical insights can be obtained regarding the role of in-waveguide attenuation in pinching-antenna design. Specifically, when $\beta$ is sufficiently small, the expression simplifies to the result in (6), which corresponds to a LoS-dominant scenario. Conversely, in moderate blockage conditions with a sufficiently large communication region (i.e., $\beta D\gg 1$), the average data rate loss $\mathbb{E}_{{\bm{\psi}}}[\Delta R]$ converges to the following asymptotic value:
@@
To further illustrate the performance advantage of pinching-antenna systems compared to the conventional fixed-antenna system, a numerical simulation of the average data rate is presented in Fig. 5. In this example, a rectangular communication area is considered with width $D_{y}=10$ m and length $D_{x}$ varying from $10$ to $80$ m, and we assume that there is a LoS channel between the pinching antenna and the user. The waveguide is deployed at the center of this region, parallel to the $x$-axis, with a fixed height $d_{v}$. For the pinching-antenna system, two schemes are considered. In the scheme with attenuation, the optimal pinching-antenna position is determined according to (5). In contrast, for the scheme without attenuation, the pinching-antenna is simply positioned directly above the user, i.e., $\tilde{x}=\bar{x}$. For the fixed-antenna system, the antenna is fixed at the geometric center of the region, i.e., $D_{x}/2,0,d_{v}]$. The simulation parameters are summarized in Table IV, which will be reused in the following simulations of this paper unless specified otherwise.
TABLE IV: Simulation parameters for Fig. 5.
@@
Refractive index ($n_{\textrm{eff}}$) $1.4$
@@
Research Insights: Single-User Pinching-Antenna Placement Optimal pinching-antenna position: When in-waveguide attenuation is small or the user is close to the waveguide, the optimal pinching-antenna position aligns with the user’s projection on the waveguide. Negligible in-waveguide attenuation in practice: For typical in-waveguide attenuation coefficients and communication region sizes, the impact of in-waveguide attenuation on system performance is often insignificant and can be safely ignored for practical system design.
@@
Recently, the authors of [41] provided new analytical insights into antenna placement for pinching-antenna-enabled multiuser communications under OMA transmission. Two practical schemes are investigated: a user-fairness-oriented (max–min) approach and a greedy-allocation-based approach. In the user-fairness-oriented scheme, the goal is to maximize the minimum achievable rate across all users by jointly optimizing the transmit power and the position of the pinching antenna. The problem is formulated as follows:
@@
where $\mathcal{M}=\{1,\ldots,M\}$ and $R^{\mathrm{OMA}}_{m},\textrm{for}~m\in{\mathcal{M}},$ denotes the data rate of user $m$. Let ${\bm{\psi}}_{m}=[x_{m},y_{m},0]$ denote the position of user $m$. Although problem (10) is non-convex, the analytical results in [41] shown that a closed-form optimal solution can be derived as
@@
NOMA: While OMA offers orthogonal user access and a simple system optimization, it can be spectrally inefficient, especially in scenarios with limited bandwidth resources. Therefore, to improve the spectral efficiency, schemes that can serve all users simultaneously are needed. In this case, NOMA presents a compelling alternative. By allowing multiple users to share the same radio resources and separating their signals in the power domain, NOMA can enhance the spectrum efficiency and the user connectivity [42, 43, 44, 45]. This approach is particularly beneficial in pinching-antenna systems where the channel conditions may vary significantly depending on user positions and antenna placement, and there is a strong motivation to serve multiple users concurrently.

### v2

@@ II-A1 System Architecture and Physical Implementation
A typical pinching-antenna system has three components: a BS acting as a central processing unit (CPU) to perform signal processing, dielectric waveguides employed as transmission media, and pinching antennas (small dielectric elements) for dynamic antenna deployment. As illustrated in Fig. 1(a), the dielectric waveguide is typically connected to the BS through a wired medium. Pinching antennas are strategically placed or dynamically activated along the waveguide, enabling on-demand radiation of electromagnetic energy into free space to form localized communication areas. This unique reconfigurability facilitates rapid deployment and flexible coverage, substantially enhancing the operational flexibility compared with conventional-antenna systems.
Figure 3: A schematic illustration of signal propagation and channel model for a SISO pinching-antenna system.
@@ II-A2 Signal Propagation and Channel Model
To illustrate the signal propagation characteristics and derive the corresponding channel model for pinching-antenna systems, we consider a basic single-input single-output (SISO) scenario. In this case, a single pinching antenna mounted on a dielectric waveguide serves one single-antenna user in the downlink, as depicted in Fig. 3. Without loss of generality, we denote the coordinates of the waveguide’s feed point as $\bm{\psi}_{0}=[0,0,d_{v}]$, where $d_{v}$ is the height of the waveguide. The locations of the pinching antenna and the user are given by $\bm{\widetilde{\psi}}^{\textrm{Pin}}=[\tilde{x},0,d_{v}]$ and $\bm{\psi}=[\bar{x},\bar{y},0]$, respectively.
@@
  • Stage 1 (In-waveguide propagation): The transmit signal $s$ is first fed into the waveguide at the feed point, after which it propagates through the waveguide’s core to the position of the pinching antenna. Due to the inherent characteristics of the waveguide, the signal experiences both attenuation and a phase shift. Specifically, the signal reaching the pinching antenna can be represented as
@@
    where $h^{\textrm{iw}}$ denotes the in-waveguide propagation channel gain (see Fig. 3), $n_{\textrm{eff}}$ represents the effective refractive index of the dielectric waveguide, $\lambda$ is the free-space wavelength, and $\alpha$ denotes the in-waveguide attenuation coefficient [^1] The in-waveguide attenuation coefficient depends on several factors, including the waveguide material, structural design, manufacturing process, and operating frequency. For instance, as reported in [32], a ceramic alumina ribbon dielectric waveguide designed for mmWave communication can exhibit an in-waveguide attenuation coefficient smaller than $0.1$ dB/m. Here, we include in-waveguide attenuation in the channel model for notational completeness; however, as will be discussed in Section III, its impact on system performance is generally negligible under mild and practically achievable system parameter settings..
@@
    where $h^{\textrm{fs}}$ denotes the free-space channel gain (see Fig. 3), and $\eta=\frac{c^{2}}{(4\pi f_{c})^{2}}$ comes from the Friis equation and captures the frequency dependence of the free-space propagation loss [34]. Here, $c$ denotes the speed of light and $f_{c}$ represents the carrier frequency.
By concatenating the channels of these two stages, the overall effective channel gain from the waveguide feed point to the user can be modeled as
@@
$=$ $\underbrace{\frac{\eta^{\frac{1}{2}}}{\|\bm{\widetilde{\psi}}^{{\rm Pin}}-\bm{\psi}\|}}_{\text{free-space path loss}}\underbrace{\exp(-\alpha\tilde{x})}_{\text{in-waveguide attenuation}}$
$\exp\bigg({-j\bigg(\underbrace{\tfrac{2\pi}{\lambda}\|\bm{\widetilde{\psi}}^{{\rm Pin}}-\bm{\psi}\|+\tfrac{2\pi}{\lambda_{g}}\tilde{x}}_{\text{overall phase shift}}\bigg)}\bigg),$ (3)
where $\lambda_{g}=\frac{\lambda}{n_{\textrm{eff}}}$ denotes the guided wavelength within the waveguide.
The channel model derived above highlights several distinctive properties of dielectric waveguide-based pinching-antenna systems. First, unlike the traditional antenna systems, where the antenna positions are fixed, the flexibility to dynamically position pinching antennas along the waveguide enables fine-grained control over path loss. By strategically adjusting antenna placement, the system can reduce propagation distances and establish strong LoS connections, thereby improving the received signal strength. Second, the overall channel gain in pinching-antenna systems captures the phase shifts resulting from both in-waveguide and free-space propagation. This dual phase accumulation, which is absent in conventional wireless channels, means that the channel phase can also be reconfigured through the mechanical repositioning of the antenna or by exploiting the waveguide’s propagation characteristics. Such tunability is not available in conventional fixed-antenna systems. Altogether, these features provide pinching-antenna systems with better adaptability in channel shaping and user-centric link creation. At the same time, they introduce new dimensions of complexity, presenting new challenges and opportunities for the development of optimization algorithms for efficient pinching-antenna system design.
@@ II-B LCX-Based Pinching Antennas
LCX, also known as radiating cable, provides a practical and user-centric means to realize flexible antenna systems, making it a natural extension of the pinching-antenna concept. As shown in Fig. 1(b), in LCX-based systems, the coaxial cable has periodic slots or apertures, which enable controlled and distributed electromagnetic energy leakage at specific locations [17, 18]. This allows the cable to function as a long, linear antenna array that supports seamless wireless coverage—an approach widely adopted in underground tunnels, subway systems, and other challenging environments [35, 36].
@@ II-C Pinching-Inspired Antenna Systems
The key idea of pinching antennas, realizing on-demand distributed radiation along conductive or engineered surfaces, can be used to design various pinching-antenna-inspired antenna array systems. In particular, signals are transmitted/received through two stages in both waveguide-based and LCX-based pinching-antenna systems. The first stage is essentially wired transmission. For example, in DOCOMO’s pinching-antenna systems, signals are passed through a dielectric waveguide. The second stage is wireless transmission. For example, in DOCOMO’s pinching-antenna system, signals are radiated by pinches and sent to the users wirelessly. Therefore, when designing pinching-inspired antenna systems, one question is whether there are other possible signal propagation media in addition to dielectric waveguides and LCX. Ideally, such media should be readily available in practical communication environments and incur a lower cost than dielectric waveguides.
One promising example for the design of pinching-inspired antenna systems is shown in Fig. 1(c), which is motivated by the following fact. In practical IoT scenarios, there are many metal wires and pipes, such as water and chimney pipes. Through surface-wave communications, these metal pipes can deliver signals to nearby users. Unlike the guided model used by dielectric-waveguide-based transmission, the signals are passed along a metal pipe in the so-called surface-wave mode, i.e., the signals are trapped and propagated on the surface of the pipe [38]. However, similar to dielectric-waveguide-based pinching antennas, the first phase of transmission is still wired propagation. Compared with dielectric-waveguide-based pinching-antenna systems, in which coverage is constrained by the length and deployment of dedicated waveguides, such pinching-inspired generalized pinching-antenna systems can exploit existing low-cost media (e.g., metal pipes and wires) and are therefore particularly attractive for extending pinching-antenna concepts to outdoor environments and providing throughput enhancement. However, this new type of pinching-antenna system based on the surface-wave mode faces two challenges. First, signals propagated via the surface-wave mode may experience higher loss and interference than signals propagated via the guided mode of dielectric waveguides. The second challenge is how to flexibly realize the second stage of pinching-antenna systems, i.e., how to dynamically radiate signals to the user. As shown in Fig. 1(c), one approach is to use active antennas instead of passive antennas, as in waveguide/LCX-based systems. Consider an IoT communication scenario in which a user uses augmented reality (AR) or virtual reality (VR) for smart factory–related training. Conventionally, the serving access point is fixed at the center of the factory and cannot provide a strong LoS link to this user. With the pinching-inspired antenna system, the user can manually attach an active antenna to the surface where a strong LoS link is available and remove the antenna after the training session.
@@
TABLE V: Comparison of representative generalized pinching-antenna systems
Aspect Dielectric-Waveguide-Based Pinching-Antenna System LCX-Based Pinching-Antenna System Pinching-Inspired Antenna System
Signal propagation medium Dielectric waveguide Coaxial cable with engineered slots Surface of conductor (e.g., metal pipes)
Propagation mode Guided mode Quasi-TEM mode Surface-wave mode
In-medium signal loss Low Medium High
Operating frequency High Low Low & High
Active antenna required? No No Yes
Deployment scenario Indoor & Semi-Outdoor Indoor & Outdoor Indoor & Outdoor
Medium cost Medium Medium Low
Reconfiguration capability High (continuous position control along waveguide) High (via slot/section configuration) High (continuous position control along metal surface)
Table V provides a comprehensive comparison of three representative generalized pinching-antenna systems, highlighting how each system leverages unique physical and signal propagation characteristics to achieve flexible, on-demand wireless coverage. The “Propagation Mode” row describes the fundamental electromagnetic mode by which signals are transported within the medium. For dielectric-waveguide-based pinching-antenna systems, signals are conveyed using the guided mode, which confines the electromagnetic wave within the dielectric structure, resulting in low in-medium signal loss and enabling high-frequency operation. In contrast, LCX-based systems utilize a quasi transverse electromagnetic (TEM) mode characteristic of coaxial cables with engineered slots; this mode provides controlled leakage for distributed coverage but incurs higher signal loss and is typically used at lower frequencies. Pinching-inspired antenna systems, built on metal pipes or radio stripes, exploit surface-wave modes in which the signals propagate along the surface of the medium and then radiate into free space at the desired position. These systems support both low- and high-frequency operation but experience higher in-medium attenuation. These systems also differ in their typical deployment scenarios. Specifically, dielectric-waveguide-based pinching-antenna systems are most suitable for indoor and semi-outdoor environments (e.g., corridors, shopping malls, factories, train stations, or stadium concourses), where waveguides can be conveniently installed and are partly shielded from harsh weather conditions. In contrast, LCX-based and pinching-inspired systems, thanks to their robust mechanical structures and compatibility with existing cables or metal infrastructure, are well suited for both indoor and outdoor deployments. In Table V, the “Reconfiguration Capability” row summarizes how the radiation points can be adjusted for each realization: dielectric-waveguide systems naturally support fine, continuous position control along the waveguide, LCX-based systems enable antenna position adjustment via slot or section configuration, and pinching-inspired systems can activate or mechanically tune individual radiating elements on existing conductors. Overall, this comparison underscores the versatility and adaptability of the generalized pinching-antenna framework, showing how diverse physical mechanisms and design choices enable tailored wireless solutions for a wide variety of coverage needs and environments.
@@ III-A Single-Waveguide Systems
Pinching-antenna systems having a single dielectric waveguide represent the most basic yet versatile architecture for channel reconfiguration in wireless networks. Depending on the number of pinching antennas deployed and activated, as well as the user configuration, the system can realize a broad range of functionalities and performance trade-offs, as shown in Fig. 4. In this subsection, we discuss two scenarios where a waveguide is equipped with a single pinching antenna and multiple pinching antennas, respectively. For each scenario, we consider both single-user and multiple-user transmissions.
@@ III-A1 Single Pinching Antenna on the Waveguide
According to [41, Lemma 1], the globally optimal solution to problem (4) can be derived in closed-form and is given by
@@
From a data rate maximization perspective, the analysis in [41] reveals that ignoring waveguide attenuation is justified, if certain mild and practically achievable conditions on the system parameters, such as waveguide height, attenuation coefficient, and the size of the communication region, are satisfied. Specifically, according to [41, Proposition 1], the average data rate loss introduced by neglecting waveguide attenuation can be approximated as
@@
Figure 5: Average rate loss $\mathbb{E}_{{\bm{\psi}}}[\Delta R]$ versus communication region side length $D$ and in-waveguide attenuation $\alpha$ for (a) $d_{v}=3$ m and (b) $d_{v}=10$ m.
@@
to ensure that the average rate loss from ignoring in-waveguide attenuation remains below a chosen threshold $\epsilon$. For example, if $d_{v}=10$ m, $\alpha=0.0092$ m-1 (corresponds to $0.08$ dB/m) [^3] For clarity, in the remainder of this paper, we report the attenuation coefficient $\alpha$ in decibels per meter (dB/m), which is commonly used in experimental studies. Its linear-scale counterpart in nepers per meter (Np/m), often used in the analytical derivations, is given by $\alpha_{\mathrm{Np/m}}=\frac{\alpha_{\mathrm{dB/m}}}{20}\ln(10)\approx 0.1151\,\alpha_{\mathrm{dB/m}}.$ For example, $\alpha=0.08$ dB/m corresponds to approximately $0.0092$ Np/m., and the target rate loss is $\epsilon=0.1$ bps/Hz, we have $D\leq 92.88$ m. This demonstrates that, for typical system parameters, the region over which the in-waveguide attenuation can be safely neglected is sufficiently large for most practical deployments.
To further illustrate the impact of $(D,\alpha)$ and different waveguide heights, Fig. 5 depicts contour plots of the average rate loss $\mathbb{E}_{{\bm{\psi}}}[\Delta R]$ for $d_{v}=3$ m and $d_{v}=10$ m. The contour lines at $0.02$, $0.05$, $0.1$, and $0.2$ bps/Hz show that, for practical attenuation levels (e.g., $\alpha\leq 0.08$ dB/m) and region sizes $D$ on the order of several tens of meters, the expected rate loss remains below about $0.05$ bps/Hz so that in–waveguide attenuation can typically be ignored, whereas only when both $\alpha$ and $D$ become simultaneously large (high-loss media and extended coverage regions) does the loss grow appreciably and attenuation-aware design become important.
More recently, the impact of in-waveguide attenuation has been studied in a more general scenario with probabilistic LoS blockage in [42]. According to the analysis in [42, Proposition 1], the average data rate loss due to ignoring in-waveguide attenuation is given by
$\mathbb{E}_{{\bm{\psi}}}[\Delta R]\approx$ $\frac{\alpha^{2}}{\beta\ln 2}\left[1\!-\!\frac{2/D}{\sqrt{\beta(1\!+\!\beta d_{v}^{2})}}\arctan\left(\frac{\sqrt{\beta}\,D/2}{\sqrt{1\!+\!\beta d_{v}^{2}}}\right)\right],\qquad{{\textrm{(bps/Hz)}}},$ (8)
where $\beta\leq 1$ is a small positive constant and denotes the blockage density parameter. It is typically chosen between $0.01$ and $0.1$ m-2 [43] and represents the density of obstacles in the propagation area. A higher $\beta$ value corresponds to a denser blockage environment, thereby reducing the probability of sustaining a LoS connection. The expression in (8) establishes how the average rate degradation is jointly influenced by key system parameters, including the in-waveguide attenuation coefficient $\alpha$, the blockage density parameter $\beta$, the waveguide deployment height $d_{v}$, and the communication side length $D$. Based on this characterization, several practical insights can be obtained regarding the role of in-waveguide attenuation in pinching-antenna design. Specifically, when $\beta$ is sufficiently small, the expression simplifies to the result in (6), which corresponds to the LoS-dominant scenario. Conversely, in moderate blockage conditions with a sufficiently large communication region (i.e., $\beta D\gg 1$), the average data rate loss $\mathbb{E}_{{\bm{\psi}}}[\Delta R]$ converges to the following asymptotic value:
@@
To further illustrate the performance advantage of pinching-antenna systems compared to the conventional fixed-antenna system, a numerical simulation of the average data rate is presented in Fig. 6. In this example, a rectangular communication area is considered with width $D_{y}=10$ m and length $D_{x}$ varying from $10$ to $80$ m, and we assume that there is a LoS channel between the pinching antenna and the user. The waveguide is deployed at the center of this region, parallel to the $x$-axis, with a fixed height $d_{v}$. For the pinching-antenna system, two schemes are considered. In the scheme with attenuation, the optimal pinching-antenna position is determined according to (5). In contrast, for the scheme without attenuation, the pinching-antenna is simply positioned directly above the user, i.e., $\tilde{x}=\bar{x}$. For the fixed-antenna system, the antenna is fixed at the center of the region, i.e., $D_{x}/2,0,d_{v}]$. The simulation parameters are summarized in Table VI, which will be reused in the following simulations of this paper unless specified otherwise.
TABLE VI: Simulation parameters for Fig. 6.
@@
Effective refractive index ($n_{\textrm{eff}}$) $1.4$
@@
Optimal pinching-antenna position: When in-waveguide attenuation is small or the user is close to the waveguide, the optimal pinching-antenna position aligns with the user’s projection on the waveguide. Negligible in-waveguide attenuation in practice: For typical in-waveguide attenuation coefficients and communication region sizes, the impact of in-waveguide attenuation on system performance is often insignificant and can be safely ignored for practical system design.
@@
Recently, the authors of [46] provided new analytical insights into antenna placement for pinching-antenna-enabled multi-user communications under OMA transmission. Two practical schemes are investigated: a user-fairness-oriented (max–min) approach and a greedy-allocation-based approach. In the user-fairness-oriented scheme, the goal is to maximize the minimum achievable rate across all users by jointly optimizing the transmit power and the position of the pinching antenna. The problem is formulated as follows:
@@
where $\mathcal{M}=\{1,\ldots,M\}$ and $R^{\mathrm{OMA}}_{m},\textrm{for}~m\in{\mathcal{M}},$ denotes the data rate of user $m$ under OMA transmission. Let ${\bm{\psi}}_{m}=[x_{m},y_{m},0]$ denote the position of user $m$. In (10c), $x_{\max}$ denotes the maximum side length along the $x$-axis. Although problem (10) is non-convex, the analytical results in [46] shown that a closed-form optimal solution can be derived as
@@
NOMA: While OMA offers orthogonal user access and a simple system optimization, it can be spectrally inefficient, especially in scenarios with limited bandwidth resources. Therefore, to improve the spectral efficiency, schemes that can serve all users simultaneously are needed. In this case, NOMA presents a compelling alternative. By allowing multiple users to share the same radio resources and separating their signals in the power domain, NOMA can enhance the spectrum efficiency and user connectivity [47, 48, 49, 50, 51]. This approach is particularly beneficial in pinching-antenna systems where the channel conditions may vary significantly depending on user positions and antenna placement, and there is a strong motivation to serve multiple users concurrently.